# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 10 Estimation of Counterfactual Evolutions
In this section, we present a general framework for counterfactual estimation based on the outcome specification in Eq. (3). Specifically, given a desired treatment allocation matrix $\bm{w}_{u}{}$, and observing $\bm{Y}{}(\bm{W}{}=\bm{w}_{o}{})$, $\bm{w}_{o}{}$, and $\bm{X}$, we aim to estimate the counterfactual evolution denoted by $\text{CFE}_{t}{(\bm{w}_{u}{})}$ as defined in (1).
To state the main theoretical result of this section, we parameterize the unknown functions in the state evolution equations. Specifically, let $g_{t}\big{(}\cdot;{\vec{\Theta}}_{g_{t}}\big{)}$ and $h_{t}{}\big{(}\cdot;{\vec{\Theta}}_{h_{t}{}}\big{)}$ represent the parameterized forms of $g_{t}(\cdot)$ and $h_{t}{}(\cdot)$ for $t=0,\ldots,T-1$, respectively. Here, $\widehat{\vec{\Theta}}_{g_{t}}$ and $\widehat{\vec{\Theta}}_{h_{t}{}}$ are vectors of appropriate dimensions, denoting the parameters of the respective functions. {assumption} Considering (27), for all $t$, we assume there exists a modification of the function $g_{t}$, which, with a slight abuse of notation, is also denoted by $g_{t}$ such that:
@@
A simple example where Assumption 10 holds is when the random variable ${M}+{M}_{t}$ is independent of all other sources of randomness. In this case, we can normalize the mean to 1 by adjusting the function $g_{t}$, scaling it by an appropriate constant factor.
For further simplicity in notation, we also define $\sigma_{t}:=\sigma+\sigma_{t}$. Then, we can collect the unknown parameters in the state evolution equations in (27) as follows:
@@
$\widehat{\tilde{\rho}}_{1}{}\leftarrow\frac{\hat{\sigma}_{0}}{N}\sum_{i=1}^{N}g_{0}\big{(}Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{0}}\big{)}^{2}$
@@
     $\widehat{H}^{i}_{t}{}\leftarrow h_{t}{}\big{(}\widehat{\tilde{\nu}}_{t}{}+\widehat{\tilde{\rho}}_{t}{}Z^{i}+\widehat{H}^{i}_{t-1}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}{}}\big{)},\;i=1,\ldots,N$
     $\widehat{\tilde{\nu}}_{t+1}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}g_{t}\big{(}\widehat{\tilde{\nu}}_{t}{}+\widehat{\tilde{\rho}}_{t}{}Z^{i}+\widehat{H}^{i}_{t-1}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{t}}\big{)}$
     $\widehat{\tilde{\rho}}_{t+1}{}\leftarrow\frac{\hat{\sigma}_{t}}{N}\sum_{i=1}^{N}g_{t}\big{(}\widehat{\tilde{\nu}}_{t}{}+\widehat{\tilde{\rho}}_{t}{}Z^{i}+\widehat{H}^{i}_{t-1}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{t}}\big{)}^{2}$
@@
[…]
In the following, we demonstrate the consistency of the results from Algorithm 3 under above assumptions.
@@
Let the conditions of Lemma 9.12 and Assumptions 10-3 hold. In particular, Assumption 9.4 holds for both observed and desired treatment allocations $\bm{w}_{o}{}$ and $\bm{w}_{u}{}$. Then, for any $t$, $\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$ provides a consistent estimator for $\nu_{t}{(\bm{w}_{u}{})}$.
@@
If the consistency in Assumption 3 is strong, i.e., $\widehat{{\cal U}}\xrightarrow{a.s.}{\cal U}$ as $N\rightarrow\infty$, then the consistency in Theorem 10.1 also holds strongly.
@@
Proof. We use an induction argument on $t\geq 1$ to prove the following more general statement. As $N\rightarrow\infty$, we show that
[…]
Then, it is straightforward to see that $\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$ also converges to $\nu_{t}{(\bm{w}_{u}{})}=\tilde{\nu}_{t}{}+\mathbb{E}\left[H_{t-1}{}\right]$ in probability, whenever $N\rightarrow\infty$.
Step 1. Let $t=1$. We begin by proving the first result in (84). To this end, we add the notation $(N)$ to the quantities associated with the system containing $N$ experimental units. We have
@@
In the last equality, we used the result of Theorem 9.8. A similar argument, combined with Assumption 9.4-(iii), yields the second result in (84) for $t=1$. The third result also follows immediately.
Fixing an arbitrary function $\psi\in{{\cal C}{\cal P}}(k)$, we also need the following intermediary result:
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{0}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\psi\big{(}Z,H_{0}{},\vec{W}_{u},{\vec{X}}\big{)}\right],$ (87)
where $\vec{W}_{u}$ represents the weak limit of $\vec{w}^{\;i}_{u}$’s and $Z\sim{\cal N}(0,1)$. To obtain this result, let $\widetilde{\psi}$ be the function such that
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{0}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},h_{0}{}\big{(}Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{0}{}}(N)\big{)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}$
$=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big{(}Z^{i},Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i},\widehat{\vec{\Theta}}_{h_{0}{}}(N)\big{)}$
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big{(}Z^{i},Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i},{\vec{\Theta}}_{h_{0}{}}\big{)}$
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[\widetilde{\psi}\big{(}Z,Y_{0}{},\vec{W}_{u},{\vec{X}},{\vec{\Theta}}_{h_{0}{}}\big{)}\right].$
Above, we used the continuous mapping theorem and Theorem 12.2 in view of Assumption 9.4. Note that $\widetilde{\psi}\in{{\cal C}{\cal P}}(k)$ because of Assumption 9.4-(ii).
@@
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\psi\big{(}Z,H_{t-1}{},\vec{W}_{u},{\vec{X}}\big{)}\right].$ (88)
[…]
$\lim_{N\rightarrow\infty}\widehat{\tilde{\nu}}_{t+1}{(N)}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}g_{t}\big{(}\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{\tilde{\rho}}_{t}{(N)}Z^{i}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{t}}(N)\big{)}$ (89)
@@
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}g_{t}\big{(}\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z^{i}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};{\vec{\Theta}}_{g_{t}}\big{)}$
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\tilde{\nu}_{t+1}{}.$
[…]
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{t}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\psi\big{(}Z,H_{t}{},\vec{W}_{u},{\vec{X}}\big{)}\right].$ (90)
For this purpose, considering $\widehat{H}^{i}_{t}{(N)}=h_{t}{}\big{(}\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{\tilde{\rho}}_{t}{(N)}Z^{i}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}{}}(N)\big{)}$, for a proper choice of the functions $\widetilde{\psi}$ and $\widetilde{\psi}^{\prime}$, we can write:
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big{(}Z^{i},\widehat{H}^{i}_{t}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big{(}Z^{i},\widehat{\tilde{\nu}}_{t}{(N)},\widehat{\tilde{\rho}}_{t}{(N)},\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i},\widehat{\vec{\Theta}}_{h_{t}{}}(N)\big{)}$
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big{(}Z^{i},\tilde{\nu}_{t}{},\tilde{\rho}_{t}{},\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i},{\vec{\Theta}}_{h_{t}{}}\big{)}$
$=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}^{\prime}\big{(}Z^{i},\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big{)}$
@@
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\widetilde{\psi}^{\prime}\big{(}Z,H_{t-1}{},\vec{W}_{u},{\vec{X}}\big{)}\right],$
where in the last line we used the induction hypothesis. Considering the definition of functions $\widetilde{\psi}$ and $\widetilde{\psi}^{\prime}$, the proof is complete. $\Box$

### v2

@@ 11 Estimation of Counterfactual Evolutions
In this section, we present a general framework for counterfactual estimation based on the outcome specification in Eq. (3) and the result of Theorem 10.9. Specifically, given a desired treatment allocation matrix $\bm{w}_{u}{}$ (corresponding to an unobserved counterfactual), and observing $\bm{Y}{}(\bm{W}{}=\bm{w}_{o}{})$, $\bm{w}_{o}{}$, and $\bm{X}$, we aim to estimate the counterfactual evolution denoted by $\text{CFE}_{t}{(\bm{w}_{u}{})}$ as defined in (1).
To state the main theoretical result of this section, we parameterize the unknown functions in the state evolution equations. Specifically, let $g_{t}\big(\cdot;{\vec{\Theta}}_{g_{t}}\big)$ and $h_{t}\big(\cdot;{\vec{\Theta}}_{h_{t}}\big)$ represent the parameterized forms of $g_{t}(\cdot)$ and $h_{t}(\cdot)$ for $t=0,\ldots,T-1$, respectively. Here, $\widehat{\vec{\Theta}}_{g_{t}}$ and $\widehat{\vec{\Theta}}_{h_{t}}$ are vectors of appropriate dimensions, denoting the unknown parameters of the respective functions. {assumption} Considering (32), for all $t$, we assume there exists a modification of the function $g_{t}$, which, with a slight abuse of notation, is also denoted by $g_{t}$ such that:
@@
An example where Assumption 11 holds is when the random variable ${M}+{M}_{t}$ is independent of all other sources of randomness. In this case, we can normalize the mean to 1 by adjusting the function $g_{t}$, scaling it by a constant factor equal to $\mathbb{E}[{M}+{M}_{t}]$.
For further simplicity in notation, we also define $\sigma_{t}:=\sqrt{\sigma^{2}+\sigma_{t}^{2}}$. Then, we can collect the unknown parameters in the state evolution equations in (32) as follows:
@@
  $\widehat{H}^{i}_{t}{}\leftarrow h_{t}\big(\widehat{\tilde{\nu}}_{t}{}+\widehat{H}^{i}_{t-1}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}}\big),\;i=1,\ldots,N$
  $\widehat{\tilde{\nu}}_{t+1}{}\leftarrow\frac{1}{N}\sum_{i=1}^{N}g_{t}\big(\widehat{\tilde{\nu}}_{t}{}+\widehat{H}^{i}_{t-1}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{t}}\big)$
@@
[Continuous parameterization] For all $t$, the mappings ${\vec{\Theta}}_{g_{t}}\mapsto g_{t}\big(\cdot;{\vec{\Theta}}_{g_{t}}\big)$ and ${\vec{\Theta}}_{h_{t}}\mapsto h_{t}\big(\cdot;{\vec{\Theta}}_{h_{t}}\big)$ are continuous functions. {assumption}[Affine outcome functions] For all $t$, the mappings $y\mapsto g_{t}(y,\cdot)$ and $y\mapsto h_{t}(y,\cdot)$ are affine; that means there exist functions $g_{t}^{1},g_{t}^{2},h_{t}^{1}$ and $h_{t}^{2}$ such that $g_{t}(y,\cdot)=yg_{t}^{1}(\cdot)+g_{t}^{2}(\cdot)$ as well as $h_{t}(y,\cdot)=yh_{t}^{1}(\cdot)+h_{t}^{2}(\cdot)$. {assumption}[Consistent parameters estimation] $\widehat{{\cal U}}$ is a consistent estimator of $\cal U$}; that is, $\widehat{{\cal U}}\xrightarrow{P}{\cal U}$ as $N\rightarrow\infty$. {assumption}[All control initialization] There is no treatment at time $t=0$; that means $W^{i}_{0}=0$, for all $i\in[N]$, and no treatment is anticipated.
In the following, we demonstrate the consistency of the results from Algorithm 4 under the above assumptions.
@@
Let the conditions of Theorem 10.9 and Assumptions 114 hold. In particular, suppose Assumption 10.4 holds for both the observed treatment allocation $\bm{w}_{o}{}$ and the target (unobserved) treatment allocation $\bm{w}_{u}{}$. Then, as $N\rightarrow\infty$, the estimator $\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$ specified in Algorithm 4 satisfies:
$\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}-\text{CFE}_{t}{(\bm{w}_{u}{})}\xrightarrow{P}0,$ (89)
for any time $t$.
@@
If the consistency in Assumption 4 holds in the strong sense, i.e., $\widehat{{\cal U}}\xrightarrow{a.s.}{\cal U}$ as $N\rightarrow\infty$, then the consistency in Theorem 11.1 is also strong. That is, the convergence in (89) holds almost surely.
@@
Proof. We use an induction argument on $t\geq 1$ to prove the following statement. As $N\rightarrow\infty$, we show that
$\widehat{\tilde{\nu}}_{t}{}\xrightarrow{P}\tilde{\nu}_{t}{},\quad\quad\quad\frac{1}{N}\sum_{i=1}^{N}\widehat{H}^{i}_{t-1}{}\xrightarrow{P}\mathbb{E}\left[H_{t-1}{}\right],$ (90)
This implies that $\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}$ also converges to $\nu_{t}{}=\tilde{\nu}_{t}{}+\mathbb{E}\left[H_{t-1}{}\right]$ in probability, whenever $N\rightarrow\infty$. Finally, applying Eq. (LABEL:eq:BSE-limit), we can write:
$\lim_{N\rightarrow\infty}\widehat{\text{CFE}}_{t}{(\bm{w}_{u}{})}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\nu_{t}{}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}Y^{i}_{t}{(\bm{w}_{u}{})}=\lim_{N\rightarrow\infty}\text{CFE}_{t}{(\bm{w}_{u}{})},$
which concludes the proof. Above, the notation $\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;$ means that the convergence holds in probability. Next, we establish the convergence results given in (90).
Step 1. Let $t=1$. We begin by proving the first result in (90). To this end, we occasionally add the notation $(N)$ to the quantities associated with the system containing $N$ experimental units to avoid ambiguity. We have
@@
In the last equality, we used the result of Theorem 10.9. The second result also follows immediately.
Fixing an arbitrary function $\psi\in{{\cal C}{\cal P}}(\frac{k}{2})$ which is also affine in its first argument, we also need the following intermediary result:
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big(\widehat{H}^{i}_{0}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\psi\big(H_{0}{},\vec{W}_{u},{\vec{X}}\big)\right],$ (93)
where $\vec{W}_{u}$ represents the weak limit of $\vec{w}^{\;i}_{u}$’s. To obtain this result, let $\widetilde{\psi}$ be the function such that
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big(\widehat{H}^{i}_{0}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big(h_{0}\big(Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{0}}(N)\big),\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)$
$=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big(Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i},\widehat{\vec{\Theta}}_{h_{0}}(N)\big)$
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}\big(Y^{i}_{0}{},\vec{w}^{\;i}_{u},{\vec{X}}^{i},{\vec{\Theta}}_{h_{0}}\big)$
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[\widetilde{\psi}\big(Y_{0}{},\vec{W}_{u},{\vec{X}},{\vec{\Theta}}_{h_{0}}\big)\right].$
Above, we used the continuous mapping theorem and Theorem 13.2 in view of Assumption 10.4. Note that $\widetilde{\psi}\in{{\cal C}{\cal P}}(\frac{k}{2})$ because of Assumption 8-(ii).
@@
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big(\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\psi\big(H_{t-1}{},\vec{W}_{u},{\vec{X}}\big)\right].$ (94)
Step 2. We show that $\widehat{\tilde{\nu}}_{t+1}{(N)}\xrightarrow{P}\tilde{\nu}_{t+1}{}$. By reusing the continuous mapping theorem, the induction hypothesis, and Assumption 4, we have
$\lim_{N\rightarrow\infty}\widehat{\tilde{\nu}}_{t+1}{(N)}$ $=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}g_{t}\big(\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{g_{t}}(N)\big)$ (95)
@@
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}g_{t}\big(\tilde{\nu}_{t}{}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};{\vec{\Theta}}_{g_{t}}\big)$
@@
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[(\tilde{\nu}_{t}{}+H_{t-1}{})g_{t}^{1}\big(\vec{W}_{u},{\vec{X}}\big)+g_{t}^{2}\big(\vec{W}_{u},{\vec{X}}\big)\right]$
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[(\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{})g_{t}^{1}\big(\vec{W}_{u},{\vec{X}}\big)+g_{t}^{2}\big(\vec{W}_{u},{\vec{X}}\big)\right]$
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[g_{t}\big(\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},\vec{W}_{u},{\vec{X}}\big)\right]=\tilde{\nu}_{t+1}{},$
where $Z_{t}$ is as given in Theorem 10.9 and in the last line we used the fact that $Z_{t}$ is independent of $\vec{W}_{u}$ and $\vec{X}$}. Similarly, one can establish $\frac{1}{N}\sum_{i=1}^{N}\widehat{H}^{i}_{t}{}\xrightarrow{P}\mathbb{E}\left[H_{t}{}\right]$. We therefore conclude the proof by demonstrating the following result:
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\psi\big(\widehat{H}^{i}_{t}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;\mathbb{E}\left[\psi\big(H_{t}{},\vec{W}_{u},{\vec{X}}\big)\right].$ (96)
For this purpose, considering $\widehat{H}^{i}_{t}{(N)}=h_{t}\big(\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}}(N)\big)$ and the fact that $\psi$ and $h_{t}$ are affine, for any $i$, we have:
$\quad\;\psi\big(\widehat{H}^{i}_{t}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)$ (97)
$=\widehat{H}^{i}_{t}{(N)}\psi^{1}\big(\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)+\psi^{2}\big(\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)$
$=h_{t}\Big(\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{H}^{i}_{t-1}{(N)},\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}}(N)\Big)\psi^{1}\big(\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)+\psi^{2}\big(\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)$
$=\left(\Big(\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{H}^{i}_{t-1}{(N)}\Big)h_{t}^{1}\Big(\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}}(N)\Big)+h_{t}^{2}\Big(\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}}(N)\Big)\right)\psi^{1}\big(\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)+\psi^{2}\big(\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big),$
where $\psi^{1}$ and $\psi^{2}$ are the affine decomposition of the $\psi$. Then, applying the continuous mapping theorem and the induction hypothesis yields the following result:
$\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\Big(\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{H}^{i}_{t-1}{(N)}\Big)h_{t}^{1}\Big(\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}}(N)\Big)\psi^{1}\big(\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)\right)$
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;$ $\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\Big(\nu_{t}{}+\widehat{H}^{i}_{t-1}{(N)}\Big)h_{t}^{1}\Big(\vec{w}^{\;i}_{u},{\vec{X}}^{i};{\vec{\Theta}}_{h_{t}}\Big)\psi^{1}\big(\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)\right)$
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;$ $\;\mathbb{E}\left[\Big(\nu_{t}{}+H_{t-1}{}\Big)h_{t}^{1}\Big(\vec{W}_{u},{\vec{X}};{\vec{\Theta}}_{h_{t}}\Big)\psi^{1}\big(\vec{W}_{u},{\vec{X}}\big)\right].$
Using Theorem 10.9 once again, we get:
$\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\Big(\widehat{\tilde{\nu}}_{t}{(N)}+\widehat{H}^{i}_{t-1}{(N)}\Big)h_{t}^{1}\Big(\vec{w}^{\;i}_{u},{\vec{X}}^{i};\widehat{\vec{\Theta}}_{h_{t}}(N)\Big)\psi^{1}\big(\vec{w}^{\;i}_{u},{\vec{X}}^{i}\big)\right)$
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize P}}}}{{=}}\;$ $\;\mathbb{E}\left[\Big(\nu_{t}{}+\rho_{t}{}Z_{t}+H_{t-1}{}\Big)h_{t}^{1}\Big(\vec{W}_{u},{\vec{X}};{\vec{\Theta}}_{h_{t}}\Big)\psi^{1}\big(\vec{W}_{u},{\vec{X}}\big)\right].$
Repeating a similar approach for other terms on the right-hand side of (97) and then reversing the steps in (97), we get the desired result, and the proof is complete. $\Box$